OpenStudy (anonymous):

In the electric current flow, it is found that the resistance (measured in the units called ohms) offered by a fixed length of wire of a given material varies inversely as the square of the diameter of the wire. If a wire 0.01 in. in diameter has a resistance of 0.306 ohm, what is the resistance of a wire of the same length and material with diameter 0.0299 in. to the nearest ten-thousandth?

OpenStudy (ranga):

R = k / d^2 (k is a constant, unknown) d1 = 0.01 and R1 = 0.306 d2 = 0.0299 R2 = ? Use the equation in line 1 for both case and divide and k will go away and you can find R2

OpenStudy (ranga):

R = k/d^2 ----- (1) 0.306 = k / (0.01)^2 ----- (2) R = k / (0.0299)^2 ----- (3) Divide (3) / (2) and solve for R